OpenStudy (anonymous):

How many different ways can a president and vice president be selected from a class of 22?

OpenStudy (anonymous):

President can be selected in 22C1 ways.. And once president is selected then there will left 21 students from which we have to select Vice president.. Now Vice president can be selected in 21C1 ways.. Just multiply them: 22C1 * 21C2 = Solve to get your answer...

OpenStudy (anonymous):

\[{^n}C_r = \frac{n!}{(n-r)! \times r!}\]